OpenStudy (anonymous):

Which is a counterexample that disproves the conjecture? A student concludes that if x is a real number, then x^2 ≤ x^4 A. 0 B. 1/4 C. 1 D. 5/4 .

OpenStudy (anonymous):

@vzfreakz

OpenStudy (anonymous):

I think 0 or 1?

OpenStudy (zzr0ck3r):

that is wrong

OpenStudy (zzr0ck3r):

0^2<= 0^4 is a true statement 1^2<= 1^4 is a true statement so these two are not good options but \((\frac{1}{4})^2=\frac{1}{16}\cancel{\le}\frac{1}{256}=(\frac{1}{4})^4\)

OpenStudy (zzr0ck3r):

you needed to find a number that makes \(x^2\le x^4\)

OpenStudy (zzr0ck3r):

a false statement